OpenStudy (maurticewilliams):

need help determine whether each equation represents a direct variation.If it does 2y=5x+1

OpenStudy (crabbyoldgamer):

If an equation is a direct variation it takes the form y=kx, where k is a constant of variation. 2y = 5x+1 y = (5/2)x + 1/2 is not of the form y = kx: this equation does not represent a direct variation.

OpenStudy (maurticewilliams):

thanks

OpenStudy (crabbyoldgamer):

You're welcome.